OpenStudy (anonymous):

evaluate the indefinite integral (sin(4x)^4)(cos(4x))dx

OpenStudy (anonymous):

I would do a \(u\) sub.

OpenStudy (anonymous):

what should I use for u?

OpenStudy (anonymous):

When you have cos/sin, you wanna sub for the one that isn't odd.

OpenStudy (anonymous):

If both are even, you have to use half angle.

OpenStudy (anonymous):

So in this case, \(u = \sin(4x)\)

OpenStudy (anonymous):

\[ \int \sin^4(4x) \frac{d(\sin(4x))}{4} \]

OpenStudy (anonymous):

so (1/20)(sin(4x)^5)

OpenStudy (anonymous):

Yeah, that looks about right.

OpenStudy (anonymous):

thank you! could you help me with a couple other questions?

OpenStudy (anonymous):

post new questions